# 00footnotetext: *Equal contribution

### v2

@@ 1 Introduction
Temporal point processes (TPPs) capture the distribution over sequences of events in time, where both the continuous arrival-times and number of events are random. They are widely used in domains such as finance, healthcare, social networks, and transportation, where understanding and forecasting event dynamics and their complex interactions is crucial. Most (neural) TPPs capture the complex interactions between events autoregressively, parameterizing a conditional intensity/density of each event given its history (Daley & Vere-Jones, 2006; Shchur et al., 2021). While natural and flexible, this factorization comes with inherent limitations: sampling scales linearly with sequence length, errors can compound in multi-step generation, and conditional generation is restricted to forecasting tasks.
@@ 2.1 Temporal Point Processes
TPPs (Daley & Vere-Jones, 2006; 2007) are stochastic processes whose realizations are finite, ordered sets of random events in time. Let ${\bm{t}}=\{t^{(i)}\}_{i=1}^{n}$, with $t^{(i)}\in[0,T]$, denote a realization of $n$ events on a bounded time interval, which can equivalently be represented by the counting process $N(t)=\sum_{i=1}^{n}\mathbf{1}\{t^{(i)}\leq t\}$ counting the number of events up to time $t$. A TPP is uniquely characterized by its conditional intensity function (Rasmussen, 2018):
@@
$p({\bm{t}})=\left(\prod_{i=1}^{n}p(t^{(i)}\mid\mathcal{H}_{t^{(i)}})\right)\left(1-F(T\mid\mathcal{H}_{t})\right)=\left(\prod_{i=1}^{n}\lambda^{*}(t^{(i)})\right)\text{ exp }\left(-\int_{0}^{T}\lambda^{*}(s)\mathrm{d}s\right),$ (2)
@@ 2.3 Flow matching with edit operations
Edit operations are encoded by mapping $({\bm{x}}_{0},{\bm{x}}_{1})$ into aligned sequences $({\bm{z}}_{0},{\bm{z}}_{1})$ in $\mathcal{Z}$, where pairs $(z_{0}^{(i)},z_{1}^{(i)})$ correspond to insertions $(\epsilon,x)$, deletions $(x,\epsilon)$, or substitutions $(x,y)$. Crucially, since the discrete flow matching dynamics in $\mathcal{Z}$ are known, they can be transferred back to $\mathcal{X}$ via $p_{s}({\bm{x}},{\bm{z}}\mid{\bm{z}}_{0},{\bm{z}}_{1})=p_{s}({\bm{z}}\mid{\bm{z}}_{0},{\bm{z}}_{1})\delta_{\operatorname{f_{rm-blanks}}({\bm{z}})}({\bm{x}})$. Then, the marginal rates $u_{s}^{{\bm{\theta}}}$ are learned in $\mathcal{X}$ by marginalizing over $\bm{z}$} with the Bregman divergence
@@ 3 Method
Let $\mathcal{T}=[0,T]$ denote the support of the TPP. We define the state space as $\mathcal{X}_{\mathcal{T}}\;=\;\bigcup_{n=0}^{\infty}\left\{(0,t^{(1)},\dots,t^{(n)},T)\in\mathcal{T}^{n}:0<t^{(1)}<\cdots<t^{(n)}<T\right\}$, denoting the set of all possible padded TPP sequences with finitely many events.
@@ 3.2 Auxiliary alignment space
We find the minimum-cost alignment between the non-boundary events of ${\bm{t}}_{0}$ and ${\bm{t}}_{1}$ with the Needleman-Wunsch algorithm (Needleman & Wunsch, 1970), i.e.,
@@
It is trivial to see from $c_{\text{sub}}$ that the aligned sequences will never encode a $\operatorname{sub}$ operation for two events that are further than $\delta$ apart. The costs for insertions and deletions and the additional condition on $c_{\text{sub}}$ ensure that the aligned sequences are jointly sorted, i.e., for any $i<j$ we have $\max\big(z_{0}^{(i)},z_{1}^{(i)}\big)<\min\big(z_{0}^{(j)},z_{1}^{(j)}\big)$ where $\min$ and $\max$ ignore $\epsilon$ tokens. This means that any interpolated ${\bm{z}}_{s}$ will be sorted. The validity of encoded $\operatorname{ins}$ and $\operatorname{del}$ operations follows immediately.
@@ 4 Experiments
We evaluate our model on seven real-world and six synthetic benchmark datasets (Omi et al., 2019; Shchur et al., 2020b; Lüdke et al., 2023; 2025). In our experiments, we compare against IFTPP (Shchur et al., 2020a), an autoregressive baseline which consistently shows state-of-the-art performance (Bosser & Taieb, 2023; Lüdke et al., 2023; Kerrigan et al., 2025). We further compare to PSDiff (Lüdke et al., 2025) and AddThin (Lüdke et al., 2023), given their strong results in both conditional and unconditional settings and their methodological similarity to our approach. All models are trained with five seeds and we select the best checkpoint based on $W_{1}$-over-$d_{\mathrm{IET}}$ against a validation set. EdiTPP, AddThin, and PSDiff are trained unconditionally but can be conditioned at inference time. [^1] To stay comparable, we employ the conditioning algorithm from Lüdke et al. (2025) for AddThin. We list the full results in Appendix C.
For forecasts, we compare predicted and target sequences by three metrics: $d_{\mathrm{Xiao}}$ introduced by Xiao et al. (2017), the mean relative error (MRE) of the event counts and $d_{\mathrm{IET}}$, which compares inter-event times to quantify the relation between events such as burstiness. In unconditional generation, we compare our generated sequences to the test set in terms of maximum mean discrepancy (MMD) (Shchur et al., 2020b) and their Wasserstein-1 distance with respect to their counts ($d_{l}$) and inter-event times ($d_{\mathrm{IET}}$). See Appendix B for details.
@@ 4.1 Unconditional generation
Table 1: Unconditional sampling performance. Bold is best, underlined second best. Ranking based on full results.
@@ 4.2 Conditional generation (Forecasting)
Table 2: Forecasting accuracy up to $T$. Bold is best, underlined second best. Ranking based on full results.
@@
$d_{\mathrm{Xiao}}$
IFTPP $6.0$ $3.9$ $6.3$ $4.7$ $2.6$ $1.8$ $3.4$
@@ 4.3 Edit efficiency
Table 3: Average number of edit operations in unconditional sampling across datasets.
@@
PSDiff 171.47 61.22 232.68
EdiTPP 132.88 31.43 28.97 193.29
The $\operatorname{sub}$ operation allows our model to modify sequences in a more targeted way when compared to PSDiff or AddThin, which have to rely on just inserts and deletes. Note that one $\operatorname{sub}$ operation can replace an insert-delete pair. Table 3 shows this results in EdiTPP using fewer edit operations than PSDiff on average even if one would count substitutions twice, as an insert and a delete. [^2] Due to its recursive definition, AddThin inserts and subsequently deletes some noise events during sampling, which results in additional edit operations compared to PSDiff. This is further amplified by the fact, that unlike EdiTPP, PSDiff and AddThin only indirectly parameterizes the transition edit rates by predicting ${\bm{t}}_{1}$ by insertion and deletion at every sampling step.
Figure 4: Changing the number of steps $k$ allows trading off compute and sample quality in terms of $d_{l}$, $d_{W_{2}}$ and $d_{\mathrm{IET}}$ at inference time.
@@ 5 Related work
The statistical modeling of TPPs has a long history (Daley & Vere-Jones, 2007; Hawkes, 1971). Classical approaches such as the Hawkes process define parametric conditional intensities, but their limited flexibility has motivated the development of neurally parameterized TPPs:
Autoregressive Neural TPP: Most neural TPPs adopt an autoregressive formulation, modeling the distribution of each event conditional on its history. These models consist of two components: a history encoder and an event decoder. Encoders are typically implemented using recurrent neural networks (Du et al., 2016; Shchur et al., 2020a) or attention mechanisms (Zhang et al., 2020a; Zuo et al., 2020; Mei et al., 2022), with attention-based models providing longer-range context at the cost of higher complexity (Shchur et al., 2021). Further, some propose to encode the history of a TPP in a continuous latent stochastic processes (Chen et al., 2020; Enguehard et al., 2020; Jia & Benson, 2019; Hasan et al., 2023). For the decoder, a wide variety of parametrizations have been explored. Conditional intensities or related measures (e.g., hazard function or conditional density), can be modeled, parametrically (Mei & Eisner, 2017; Zuo et al., 2020; Zhang et al., 2020a), via neural networks (Omi et al., 2019), mixtures of kernels (Okawa et al., 2019; Soen et al., 2021; Zhang et al., 2020b) and mixture distributions (Shchur et al., 2020a). Generative approaches further enhance flexibility: normalizing flow-based (Shchur et al., 2020b), GAN-based (Xiao et al., 2017), VAE-based (Li et al., 2018), and diffusion-based decoders (Lin et al., 2022; Yuan et al., 2023) have all been proposed. While expressive, autoregressive TPPs are inherently sequential, which makes sampling scale at least linearly with sequence length, can lead to error accumulation in multi-step forecasting and limit conditional generation to forecasting.
@@ Appendix A Model Parameters
All MLPs have input and output sizes of $H$, except for the final MLP whose output size is determined by the number of $\lambda$ and $Q$ parameters of the rate. The MLPs have a single hidden layer of size $4H$. The sinusoidal embeddings map a scalar $s\in[0,1]$ to a vector of length $H$. In contrast to Havasi et al. (2025), we choose a cosine $\kappa$ schedule $\kappa(s)=1-{\cos\big(\frac{\pi}{2}s\big)}^{2}$ as proposed by Nichol & Dhariwal (2021) for diffusion models as it improved results slightly compared $\kappa(s)=s^{3}$.
@@
Appendix B Metrics

### v3

@@ 1 Introduction
Temporal point processes (TPPs) capture the distribution over sequences of events in time, where both the continuous arrival-times and number of events are random. They are widely used in domains such as finance, healthcare, social networks, and transportation, where understanding and forecasting event dynamics and their complex interactions is crucial. Most (neural) TPPs capture the complex interactions between events autoregressively, parameterizing a conditional intensity/density of each event given its history (Daley and Vere-Jones, 2006; Shchur et al., 2021). While natural and flexible, this factorization comes with inherent limitations: sampling scales linearly with sequence length, errors can compound in multi-step generation, and conditional generation is restricted to forecasting tasks.
@@ 2.1 Temporal Point Processes
TPPs (Daley and Vere-Jones, 2006; 2007) are stochastic processes whose realizations are almost surely finite, ordered sets of random events in time. Let ${\bm{t}}=\{t^{(i)}\}_{i=1}^{n}$, with $t^{(i)}\in[0,T]$, denote a realization of $n$ events on a bounded time interval, which can equivalently be represented by the counting process $N(t)=\sum_{i=1}^{n}\mathbf{1}\{t^{(i)}\leq t\}$ counting the number of events up to time $t$. A TPP is uniquely characterized by its conditional intensity function (Rasmussen, 2018):
@@
$p({\bm{t}})=\left(\prod_{i=1}^{n}p(t^{(i)}\mid\mathcal{H}_{t^{(i)}})\right)\left(1-F(T\mid\mathcal{H}_{t^{(i)}})\right)=\left(\prod_{i=1}^{n}\lambda^{*}(t^{(i)})\right)\text{ exp }\left(-\int_{0}^{T}\lambda^{*}(s)\mathrm{d}s\right),$ (2)
@@ 2.3 Flow matching with edit operations
Edit operations are encoded by introducing a blank token $\epsilon$ and mapping $({\bm{x}}_{0},{\bm{x}}_{1})$ into aligned sequences $({\bm{z}}_{0},{\bm{z}}_{1})$ in $\mathcal{Z}$, where pairs $(z_{0}^{(i)},z_{1}^{(i)})$ correspond to insertions $(\epsilon,x)$, deletions $(x,\epsilon)$, or substitutions $(x,y)$. Crucially, since the discrete flow matching dynamics in $\mathcal{Z}$ are known, they can be transferred back to $\mathcal{X}$ via $p_{s}({\bm{x}},{\bm{z}}\mid{\bm{z}}_{0},{\bm{z}}_{1})=p_{s}({\bm{z}}\mid{\bm{z}}_{0},{\bm{z}}_{1})\delta_{\operatorname{f_{rm-blanks}}({\bm{z}})}({\bm{x}})$, by removing $\epsilon$’s with $\operatorname{f_{rm-blanks}}$. Then, the marginal rates $u_{s}^{{\bm{\theta}}}$ are learned in $\mathcal{X}$ by marginalizing over $\bm{z}$} with the Bregman divergence
@@ 3 Method
Let $\mathcal{T}=[0,T]$ denote the support of the TPP. We define the state space as $\mathcal{X}_{\mathcal{T}}\;=\;\bigcup_{n=0}^{\infty}\left\{(0,t^{(1)},\dots,t^{(n)},T)\in\mathcal{T}^{n}:0<t^{(1)}<\cdots<t^{(n)}<T\right\}$, denoting the set of all possible padded TPP sequences with finitely many events. Note that the padding values are introduced for notational simplicity when defining the edit operations on $\mathcal{T}$.
@@ 3.1 Edit operations
Our choice of $\operatorname{ins}$, $\operatorname{sub}$ and $\operatorname{del}$ ensures three key properties: (i) the resulting event sequences remain valid TPPs, (ii) the number of valid operations, e.g. $\operatorname{ins}({\bm{t}},i,j)$, is independent of the position $i$, which is necessary for efficient parameterization, and (iii) at most one unique operation can transition between any two states, which significantly reduces the complexity of the training loss in Section 3.3. While these properties are comparably simple to achieve for token sequences in language modeling (Havasi et al., 2025), where any token can replace any other, they require special care in the case of TPPs. $\operatorname{del}$ and $\operatorname{sub}$ are defined to ensure that the resulting event sequence remains in increasing order and that the padding events $t^{(0)}$ and $t^{(n+1)}$ remain in place. $\operatorname{del}$ transitions are unique because the removed event determines exactly which deletion occurred. Similarly, $\operatorname{sub}$ transitions are unique because the original position of the substituted event disambiguates the operation, even though two distinct $\operatorname{sub}$ operations may yield the same substituted event value. To achieve uniqueness for $\operatorname{ins}$, the insertion bins corresponding to $\operatorname{ins}({\bm{t}},i,j)$ have to be mutually disjoint for any $i,j$ since insertions lack a removed event to disambiguate them. We achieve this by sizing the bins relative to the distance between $t^{(i)}$ and $t^{(i+1)}$.
@@ 3.2 Auxiliary alignment space
We find the minimum-cost alignment between the non-boundary events of ${\bm{t}}_{0}$ and ${\bm{t}}_{1}$ with the Needleman-Wunsch algorithm (Needleman and Wunsch, 1970), i.e.,
@@
We carefully craft the cost functions in Eq. 13, to guarantee that the minimum-cost alignment corresponds to $\operatorname{ins}$, $\operatorname{sub}$ and $\operatorname{del}$ operations as we define them in Section 3.1. With the $|t_{0}^{(i)}-t_{1}^{(j)}|<\delta$ condition in $c_{\text{sub}}$, we ensure that the aligned sequences will never encode a $\operatorname{sub}$ operation for two events that are further than $\delta$ apart. The costs for insertions and deletions and the additional condition on $c_{\text{sub}}$ ensure that the aligned sequences are jointly sorted, i.e., for any $i<j$ we have $\max\big(z_{0}^{(i)},z_{1}^{(i)}\big)<\min\big(z_{0}^{(j)},z_{1}^{(j)}\big)$ where $\min$ and $\max$ ignore $\epsilon$ tokens. This means that any interpolated ${\bm{z}}_{s}$ is sorted by construction. The validity of encoded $\operatorname{ins}$ and $\operatorname{del}$ operations follows immediately.
@@ 4 Experiments
We evaluate our model on seven real-world and six synthetic benchmark datasets (Omi et al., 2019; Shchur et al., 2020b; Lüdke et al., 2023; 2025). In our experiments, we compare against IFTPP (Shchur et al., 2020a), an autoregressive baseline which consistently shows state-of-the-art performance (Bosser and Taieb, 2023; Lüdke et al., 2023; Kerrigan et al., 2025). We further compare to PSDiff (Lüdke et al., 2025) and AddThin (Lüdke et al., 2023), given their strong results in both conditional and unconditional settings and their methodological similarity to our approach. All models are trained with five seeds and we select the best checkpoint based on $W_{1}$-over-$d_{\mathrm{IET}}$ against a validation set. EdiTPP, AddThin, and PSDiff are trained unconditionally but can be conditioned at inference time. [^1] To stay comparable, we employ the conditioning algorithm from Lüdke et al. (2025) for AddThin. We list the full results in Section E.3.
For forecasts, we compare predicted and target sequences by three metrics: $d_{\mathrm{Xiao}}$ introduced by Xiao et al. (2017), the mean relative error (MRE) of the event counts and $d_{\mathrm{IET}}$, which compares inter-event times to quantify the relation between events such as burstiness. In unconditional generation, we compare our generated sequences to the test set in terms of maximum mean discrepancy (MMD) (Shchur et al., 2020b) and their Wasserstein-1 distance with respect to their counts ($d_{l}$) and inter-event times ($d_{\mathrm{IET}}$). See Appendix C for details.
@@ 4.1 Unconditional generation
Table 1: Unconditional sampling performance. Bold is best, underlined second best. Ranking follows full results in Section E.3 and results are grouped if they fall within the std of the best member.
@@ 4.2 Conditional generation (Forecasting)
Table 2: Forecasting accuracy up to $T$. Bold is best, underlined second best. Ranking follows full results in Section E.3 and results are grouped if they fall within the std of the best member.
@@
$d_{\mathrm{Xiao}}$ IFTPP $6.0$ $3.9$ $6.3$ $4.7$ $2.6$ $1.8$ $3.4$
@@ 4.3 Edit efficiency
The $\operatorname{sub}$ operation allows our model to modify sequences in a more targeted way when compared to PSDiff or AddThin, which have to rely on just inserts and deletes.
Table 3: Average number of edit operations in unconditional sampling across datasets. Full result in Table 13.
@@
PSDiff 173.48 61.04 0.00 234.52
EdiTPP 137.42 33.08 29.16 199.65
Note that one $\operatorname{sub}$ operation can replace an insert-delete pair. Table 3 shows this results in EdiTPP using fewer edit operations than PSDiff on average even if one would count substitutions twice, as an insert and a delete. [^2] Due to its recursive definition, AddThin inserts and subsequently deletes some noise events during sampling, which results in additional edit operations compared to PSDiff. This is further amplified by the fact, that unlike EdiTPP, PSDiff and AddThin only indirectly parameterizes the transition edit rates by predicting ${\bm{t}}_{1}$ by insertion and deletion at every sampling step.
Figure 4: Changing the number of steps $k$ allows trading off compute and sample quality in terms of $d_{l}$, $d_{\mathrm{IET}}$ and MMD at inference time.
@@ 5 Related work
The statistical modeling of TPPs has a long history (Daley and Vere-Jones, 2007; Hawkes, 1971). Classical approaches such as the Hawkes process define parametric conditional intensities, but their limited flexibility has motivated the development of neurally parameterized TPPs:
Autoregressive Neural TPP: Most neural TPPs adopt an autoregressive formulation, modeling the distribution of each event conditional on its history. These models consist of two components: a history encoder and an event decoder. Encoders are typically implemented using recurrent neural networks (Du et al., 2016; Shchur et al., 2020a) or attention mechanisms (Zhang et al., 2020a; Zuo et al., 2020; Mei et al., 2022), with attention-based models providing longer-range context at the cost of higher complexity (Shchur et al., 2021). Further, some propose to encode the history of a TPP in a continuous latent stochastic processes (Chen et al., 2020; Enguehard et al., 2020; Jia and Benson, 2019; Hasan et al., 2023). For the decoder, a wide variety of parametrizations have been explored. Conditional intensities or related measures (e.g., hazard function or conditional density), can be modeled, parametrically (Mei and Eisner, 2017; Zuo et al., 2020; Zhang et al., 2020a), via neural networks (Omi et al., 2019), mixtures of kernels (Okawa et al., 2019; Soen et al., 2021; Zhang et al., 2020b) and mixture distributions (Shchur et al., 2020a). Generative approaches further enhance flexibility: normalizing flow-based (Shchur et al., 2020b), GAN-based (Xiao et al., 2017), VAE-based (Li et al., 2018), and diffusion-based decoders (Lin et al., 2022; Yuan et al., 2023) have all been proposed. While expressive, autoregressive TPPs are inherently sequential, which makes sampling scale at least linearly with sequence length, can lead to error accumulation in multi-step forecasting and limit conditional generation to forecasting.
@@ 6 Conclusion
Acknowledgments
We want to thank the Munich Center for Machine Learning for providing compute resources.
@@ Appendix A Model Parameters
All MLPs have input and output sizes of $H$, except for the final MLP whose output size is determined by the number of $\lambda$ and $Q$ parameters of the rate. The MLPs have a single hidden layer of size $4H$. The sinusoidal embeddings map a scalar $s\in[0,1]$ to a vector of length $H$. In contrast to Havasi et al. (2025), we choose a cosine $\kappa$ schedule $\kappa(s)=1-{\cos\big(\frac{\pi}{2}s\big)}^{2}$ as proposed by Nichol and Dhariwal (2021) for diffusion models as it improved results slightly compared $\kappa(s)=s^{3}$.
@@
Appendix B Data
B.1 Synthetic Datasets
The six synthetic datasets were generated by Shchur et al. (2020b) following the simulation procedures detailed in Section 4.1 of Omi et al. (2019). Each dataset contains 1,000 sequences supported on the interval $T=[0,100]$. They cover a diverse set of temporal dynamics, defined as follows:
Hawkes Processes (H1, H2).
Hawkes processes capture self-exciting features of temporal point processes. The two Hawkes processes are parameterized as follows:
$\lambda(t\mid\mathcal{H}_{t})=\mu\;+\;\sum_{t_{i}<t}\sum_{j=1}^{M}\alpha_{j}\beta_{j}\exp\{-\beta_{j}(t-t_{i})\},$
with H1 ($M=1$, $\mu=0.2$, $\alpha_{1}=0.8$, $\beta_{1}=1.0$) and H2 ($M=2$, $\mu=0.2$, $\alpha_{1}=0.4$, $\beta_{1}=1.0$, $\alpha_{2}=0.4$, $\beta_{2}=20.0$).
Non-stationary Poisson Process (NSP).
A periodic time-varying intensity:
$\lambda(t\mid\mathcal{H}_{t})=0.99\sin\!\left(\frac{2\pi t}{20000}\right)+1.$
Stationary Renewal Process (SR).
Inter-event times $\tau_{i}=t_{i+1}-t_{i}$ are i.i.d. from a log-normal distribution (mean $1.0$, std. $6.0$): this produces bursty patterns with short activity bursts followed by long silent periods.
Non-stationary Renewal Process (NSR).
A stationary renewal process is first generated using a gamma distribution (mean $1.0$, std. $0.5$), then timestamps are time-warped by
$t^{\prime}_{i}=\int_{0}^{t_{i}}r(s)\,ds,\qquad r(t)=0.99\sin\!\left(\frac{2\pi t}{20000}\right)+1.$
This induces temporally varying expected inter-event intervals while preserving local correlations.
Self-correcting Process (SC).
The intensity grows with the time elapsed since the last event:
$\lambda(t\mid\mathcal{H}_{t})=\exp\!\left(t-\sum_{t_{i}<t}1\right).$
This discourages extended silent periods and promotes regular spacing.
B.2 Real-World Datasets
We use the seven real-world datasets proposed by (Shchur et al., 2020b):
PG (PUBG) represents death-event timestamps from matches of PUBG. R/C (Reddit-Comments) consists of comment timestamps within the first 24 hours of threads posted on r/askscience, covering 01.01.2018–31.12.2019. R/P (Reddit-Submissions) captures daily submission timestamps from r/politics, covering 01.01.2017–31.12.2019. Tx (Taxi) are taxi pick-up events in the southern part of Manhattan, New York. Tw (Twitter) covers tweet timestamps of user ID 25073877, collected over multiple years. Y/A (Yelp-Airport) consists of check-in events at McCarran International Airport (27 users, year 2018). Lastly, Y/M (Yelp-Mississauga) presents check-ins for businesses in the city of Mississauga (27 users, year 2018).
Table 6: Summary statistics for all synthetic and real-world datasets. $\bm{\tau}$ is the average inter-event time.
Full Name Abbrev. # Seq. Mean Len. Support [0, T] $\bm{\tau}$
Hawkes 1 H1 1000 95.4 100 $1.01\pm 2.38$
Hawkes 2 H2 1000 97.2 100 $0.98\pm 2.56$
Nonstationary Poisson NSP 1000 100.3 100 $0.99\pm 2.22$
Nonstationary Renewal NSR 1000 98.0 100 $0.98\pm 1.83$
Self-Correcting SC 1000 100.2 100 $0.99\pm 0.71$
Stationary Renewal SR 1000 109.2 100 $0.83\pm 2.76$
PUBG PG 3001 76.5 38 minutes $0.41\pm 0.56$
Reddit Comments R/C 1356 295.7 24 hours $0.07\pm 0.28$
Reddit Submissions R/P 1094 1129.0 24 hours $0.02\pm 0.03$
Taxi Pick-ups (Manhattan) Tx 182 98.4 24 hours $0.24\pm 0.40$
Twitter Activity Tw 2019 14.9 24 hours $1.26\pm 2.80$
Yelp Check-ins (Airport) Y/A 319 30.5 24 hours $0.77\pm 1.10$
Yelp Check-ins (Mississauga) Y/M 319 55.2 24 hours $0.43\pm 0.96$
Appendix C Metrics